IDENTIFICATION OF NEW CIS VIBRATIONAL LEVELS IN THE S_{1} STATE OF C_{2}H_{2}

\underline{J. H. BARABAN}, P. B. CHANGALA, R. G. SHAVER, R. W. FIELD, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139, USA; J. F. STANTON, Institute for Theoretical Chemistry, Departments of Chemistry and Biochemistry, The University of Texas at Austin, Austin, Texas 78712; A. J. MERER, Institute of Atomic and Molecular Sciences, Academia Sinica, Taipei 10617, Taiwan.

Although the S_{1} (Ã ^{1}A_{u}) state of the trans conformer of acetylene has been known for many years, the corresponding S_{1} (Ã ^{1}A_{2}) state of the cis conformer was only discovered recently.^{a} Transitions to it from the ground state are electronically forbidden, but its vibrational levels acquire intensity by tunneling through the isomerization barrier and interacting with levels of the trans conformer. We have recently identified two new vibrational levels (3^{2} and 4^{1}6^{1}) of the cis conformer of S_{1} C_{2}H_{2}, bringing the total number of levels observed to six out of an expected ten up to the energies studied in this work. The appearance of these levels in IR-UV double resonance LIF spectra will be discussed, along with their vibrational assignments. Experimentally determined vibrational parameters and ab initio anharmonic force fields for both the trans and cis conformers will be presented as part of the evidence supporting these assignments. These results shed new light on the vibrational level structure of both conformers in this isomerizing system.

aA. J. Merer, A. H. Steeves, J. H. Baraban, H. A. Bechtel, and R. W. Field. J. Chem. Phys., 134(24):244310, 2011.